# InterFormer: Towards Effective Heterogeneous Interaction Learning for Click-Through Rate Prediction

### v2

@@ Abstract.
Click-through rate (CTR) prediction, which predicts the probability of a user clicking an ad, is a fundamental task in recommender systems. The emergence of heterogeneous information, such as user profile and behavior sequences, depicts user interests from different aspects. A mutually beneficial integration of heterogeneous information is the cornerstone towards the success of CTR prediction. However, most of the existing methods suffer from two fundamental limitations, including (1) insufficient inter-mode interaction due to the unidirectional information flow between modes, and (2) aggressive information aggregation caused by early summarization, resulting in excessive information loss. To address the above limitations, we propose a novel module named InterFormer to learn heterogeneous information interaction in an interleaving style. To achieve better interaction learning, InterFormer enables bidirectional information flow for mutually beneficial learning across different modes. To avoid aggressive information aggregation, we retain complete information in each data mode and use a separate bridging arch for effective information selection and summarization. Our proposed InterFormer achieves state-of-the-art performance on three public datasets and a large-scale industrial dataset.
@@ 1. Introduction
Click-through rate (CTR) prediction, which aims to predict the probability of a user clicking an ad or item, is the fundamental task for various applications such as online advertising and recommender systems (Song et al., 2019; Lyu et al., 2020; Zhang et al., 2022, 2024). The quality of CTR prediction significantly influences the company revenue and user experience, drawing extensive attention from both academia and industry (Zhou et al., 2018, 2019). For example, in online ad bidding, accurate CTR prediction helps advertisers optimize their bids and target the most receptive audiences. In content recommendation, it enables platforms to suggest more relevant content to users.
@@
To achieve better CTR prediction, it is crucial to capture the user interests in the evolving environment (Zhou et al., 2018; Lyu et al., 2020; Wang et al., 2019). The abundance of heterogeneous information presents both opportunities and challenges. On the one hand, heterogeneous information depicts user interests from different aspects, providing diverse context (Zhang et al., 2017). For instance, global information, e.g., user profile and context features, offers a static view on general user interests, while user behavior sequences provide substantial information for modeling dynamic user interests (Wang et al., 2019). On the other hand, the heterogeneous nature of the data requires different modeling approaches and careful integration across different information modes (Zhang et al., 2017). For example, while modeling interactions among global information is critical to personalized recommendation (Rendle, 2010; Lian et al., 2018; Wang et al., 2021b), capturing sequential dependencies is the major focus for user behavior modeling (Sun et al., 2019; Chen et al., 2019).
@@
Most of the existing CTR prediction models fall into two categories, including non-sequential models and sequential models. Non-sequential models focus on learning informative embeddings through feature interaction via inner-product (Lian et al., 2018; Sun et al., 2021), MLP (Wang et al., 2017, 2021b) and deep structured semantic model (Huang et al., 2013; Elkahky et al., 2015), but ignore the sequential information in user behaviors. Sequential models, in contrast, employ additional modules like CNN (Tang and Wang, 2018), RNN (Sun et al., 2019; Zhou et al., 2018) and Attention modules (Lyu et al., 2020; Zhou et al., 2019; Zhai et al., 2024), to capture the sequential dependencies in user behaviors. Promising as it might be, existing sequential methods mostly employ a unidirectional information flow, where global information is used to guide sequence learning, while the reverse information flow from sequence to global information is largely ignored, resulting in insufficient inter-mode interaction. For example, global information often captures long-term interests, while sequence information reveals momentary interests, such as a sudden focus on a specific category of products, which can enhance the global context with immediate preference. Besides, due to the computational challenges of performing interaction learning among numerous global features and lengthy sequences, aggressive feature aggregation, e.g., sequence summation (Zhou et al., 2018), pooling (Xiao et al., 2020), and concatenation (Zhou et al., 2019), is often performed in the early stages, inevitably leading to excessive information loss.
@@
In light of the above limitations, we propose a novel heterogeneous interaction learning module named InterFormer, whose ideas are two-fold. To avoid insufficient inter-mode interaction, we enable bidirectional information flows between different modes, such that global and sequence learning are performed in an interleaving style. Specifically, to learn context-aware sequence embeddings, global sumamrization guides sequence learning via Personalized FeedForward Network (PFFN) and Multihead Attention (MHA) (Vaswani et al., 2017). To learn behavior-aware global embeddings, sequence summarization instructs global learning via an interaction module. To mitigate aggressive information aggregation, we adopt MHA for effective information selection, based on which the one-to-one mappings between input and output tokens are retained till the final interaction. Note that our framework is compatible with various interaction learning models like DCNv2 (Wang et al., 2021b), DHEN (Zhang et al., 2022), etc.
@@
  • Model Design. We propose a novel heterogeneous interaction learning framework named InterFormer for effective feature interaction and selective information aggregation. To our best knowledge, the proposed InterFormer is the first model to address the mutual benefits in heterogeneous interaction learning.
  • Experiments and Analysis. We carry out extensive experiments on the proposed InterFormer with up to 0.14% AUC improvement on benchmark datasets and 0.15% Normalized Entropy (NE) gain on internal large-scale dataset. In addition, we have conducted studies on internal deployment at Meta and find that InterFormer exhibits promising scaling results in both feature scaling and model scaling.
@@ 2.2. Sequential Methods
With the recent emergence of sequence information, e.g., user interaction history, in recommender systems, extensive efforts have been made to capture the evolving user interest. A key challenge behind sequential methods is to combine the sequential information with the heterogeneous global information in a mutually beneficial manner. Markov models (Shani et al., 2005; He and McAuley, 2016a; Yang et al., 2020) consider sequential data as a stochastic process over discrete random variables, but the oversimplified Markovian property limits the model capability in capturing long-term sequential dependencies (Quadrana et al., 2018). To model the long-term dependencies, RNN and Attention mechanism are employed as the backbone module for many sequential methods. For example, to extract the evolving user interest, various attention-based networks are designed (Zhou et al., 2019; Lyu et al., 2020), and Transformer architecture (Devlin et al., 2018; Vaswani et al., 2017) is adopted for sequential modeling (Sun et al., 2019; Chen et al., 2019). Besides, to model multi-faceted user interests, (Xiao et al., 2020; Han et al., 2024) propose to capture multiple user interests from multi-behavior sequences. More recently, TransAct (Xia et al., 2023) adopt a hybrid ranking model to combine real-time user actions for immediate preference and batch user representations for long-term interests. LiRank (Borisyuk et al., 2024) improves CTR prediction at LinkedIn by ensembling multiple interaction modules, accelerated by quantization and vocabulary compression. CARL (Chen et al., 2024) achieves fast inference on large-scale recommendation in Kuaishou by utilizing cached results, monitored by a reinforcement learning-based framework. While most of the existing sequential methods leverage global information for personalized sequence modeling, sequence information is rarely explored for better global learning. We believe such unidirectional design limits the expressiveness of the learned embeddings, and a bidirectional information flow between different data modes is the key towards better heterogeneous interaction learning.
@@ 3. Preliminaries
$\mathbf{x}_{j}^{(l)}$ $j$-th global feature of layer $l$
$\mathbf{s}_{t}^{(l)}$ sequence feature of layer $l$ at timestamp $t$
@@
Table 1 summarizes the main symbols and notations used throughout the paper. We use bold uppercase letters for matrices (e.g., $\mathbf{X}$), bold lowercase letters for vectors (e.g., $\mathbf{x}$), and lowercase letters for scalars (e.g., $n$). The element at the $i$-th row and $j$-th column of a matrix $\mathbf{X}$ is denoted as $\mathbf{X}(i,j)$. The transpose of $\mathbf{X}$ is denoted by the superscript ${\scriptscriptstyle\mathsf{T}}$ (e.g., $\mathbf{X}^{\scriptscriptstyle\mathsf{T}}$). We use superscript $u$ to denote users, and subscripts $i$ and $t$ to denote item and timestamp, respectively (e.g., $y^{u}_{i_{t}}$). We use $\mathbf{x}_{j}^{(l)}$ to denote the $j$-th global feature of the $l$-th layer, and $\mathbf{s}_{t}^{(l)}$ to denote the sequence feature of the $l$-th layer at timestamp $t$. We consider the scenario with $m$ dense features, $n$ sparse features, and $k$ sequence features of length $T$. We use $d$ to denote the embedding dimension of the CTR model.
@@ 3.1. Click-Through Rate (CTR) Prediction
CTR prediction aims to estimate the probability of a user clicking on an item given heterogeneous information such as global context and behavior sequences. Formally, given a user set $\mathcal{U}$ and an item set $\mathcal{I}$, the interaction sequence of a user $u\in\mathcal{U}$ is defined as $S^{u}=[i^{u}_{1},i^{u}_{2},\dots,i^{u}_{T}]$, where $i^{u}_{t}\in\mathcal{I}$ is the item interacted at time step $t$ by user $u$. CTR prediction estimates the probability of a user clicking on a new item $i^{u}_{T+1}$, denoted as $y^{u}_{i_{T+1}}$, given the historical interaction sequence $S^{u}$. Formally, we seek to learn a function $f:\mathcal{U}\times\mathcal{I}\times\mathcal{S}\rightarrow[0,1]$, where $\mathcal{S}$ is the set of all possible sequences, such that:
@@ 3.2. Feature Interaction
By stacking multiple cross layers, DCNv2 is capable of modeling explicit high-order interactions. Besides, a deep network, which is essentially an MLP, is employed to model implicit interactions between features, and the final prediction is further obtained by combining the outputs of both networks.
Deep Hierarchical Ensemble Network (DHEN) (Zhang et al., 2022). DHEN leverages the strengths of heterogeneous interaction modules and learns a hierarchy of interactions of different orders. The key idea behind DHEN is to use a layered structure where each layer is an ensemble of multiple heterogeneous interaction modules.
Formally, given the concatenation of $m$ input features $\mathbf{X}^{(l)}\in\mathbb{R}^{d\times m}$, the output of the $l$-th DHEN layer is computed as:
@@
By ensembling different modules (e.g., inner product, DCNv2, etc.), DHEN leverages their complementary strengths within a unified framework. Besides, by stacking multiple DHEN layers, high-order feature interactions can be captured.
@@ 3.3. Attention Mechanism
Attention mechanisms (Bahdanau et al., 2014; Vaswani et al., 2017) have become an integral part of sequence modeling, serving as the core component behind various model designs, such as Transformer (Vaswani et al., 2017) and BERT (Devlin et al., 2018). In this section, we briefly review attention mechanisms, including Multi-Head Attention and Pooling by Multihead Attention (PMA).
@@
(2) $\vspace{-2pt}\text{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\mathbf{QK}^{\scriptscriptstyle\mathsf{T}}/\sqrt{d_{k}}\right)\mathbf{V},\vspace{-2pt}$
where $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{n\times d_{k}}$ are query, key and value representations derived from the input sequence $\mathbf{S}$ via linear projections $\mathbf{Q}=\mathbf{S}^{\scriptscriptstyle\mathsf{T}}\mathbf{W}^{Q},\mathbf{K}=\mathbf{S}^{\scriptscriptstyle\mathsf{T}}\mathbf{W}^{K},\mathbf{V}=\mathbf{S}^{\scriptscriptstyle\mathsf{T}}\mathbf{W}^{V}$.
@@
(3) $\vspace{-2pt}\text{MHA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\left[\textbf{head}_{1}\|\ldots\|\textbf{head}_{h}\right]\mathbf{W}^{O},\vspace{-2pt}$
@@
(4) $\vspace{-2pt}\text{PMA}(\mathbf{Q}_{\text{PMA}},\mathbf{S})=\text{MHA}(\mathbf{Q}_{\text{PMA}},\mathbf{K},\mathbf{V}).\vspace{-2pt}$
@@ 4. Methodology
Figure 1. An overview of the InterFormer model architecture. Each InterFormer block consists of three parts, including (1) Global Arch (orange) to learn behavior-aware global embeddings given sequence queries; (2) Sequence Arch (blue) to learn context-aware sequence embeddings given global queries; (3) Bridging Arch (green) to exchange information between global and sequence arch. Note that the dashed green line for CLS token appending only happens at the first layer.
In this section, we present our proposed InterFormer. We first introduce the preprocessing module in Section 4.1, followed by three major modules, including Global Arch, Sequence Arch and Bridging Arch. To learn behavior-aware global embeddings, the Global Arch is proposed to model feature interactions (Section 4.2). To learn context-aware sequence embeddings, the Sequence Arch is proposed to optimize sequence (Section 4.3). The Bridging Arch connects Global and Sequence Archs, enabling effective information summarization and exchange between different modes (Section 4.4). An overview of the InterFormer architecture is shown in Figure 1.
@@ 4.1. Feature Preprocessing
Global Feature Preproessing. We consider two types of global features, including dense features like user age and item price, and sparse features like user id and item category. To unify the heterogeneity in global features, the global preprocessing module transforms features into embeddings of the same dimensionality.
@@
By concatenating the dense and sparse embedding vectors, the input global embedding matrix can be obtained as follows
@@
Sequence Feature Preprocessing. Similar to the global feature preprocessing, an embedding layer is employed such that each interacted item in the sequence is mapped to a $d$-dimensional vector $\mathbf{s}_{t}$, and the user behavior sequences is represented as the concatenation of item embeddings, i.e., $\mathbf{S}^{(0)}=\left[\mathbf{s}_{1}^{(0)}\|\dots\|\mathbf{s}_{T}^{(0)}\right]\in\mathbb{R}^{d\times T}$.
Real-world scenarios often encounter multiple sequences originating from different user actions (e.g., click and conversion)or various platforms. Moreover, due to the uncertainty of user behaviors, these sequences often contain noisy and irrelevant user-item interactions. To address this, a MaskNet (Wang et al., 2021a) is employed to unify multiple sequences and filter out the internal noises via self-masking.
@@
where $\text{MLP}_{\text{mask}}:\mathbb{R}^{kd\times T}\to\mathbb{R}^{kd\times T}$ generates the self-masking to select relevant information from the input sequence, and $\text{MLP}_{\text{lce}}:\mathbb{R}^{kd\times T}\to\mathbb{R}^{d\times T}$ linearly combines multiple sequences into one, matching the dimensionalities of global and sequence features.
4.2. Global Arch: Towards Behavior-Aware Interaction Learning
Global features, such as user profile and ad content, provide substantial information in understanding user preference over a specific item. Modeling the interaction among global features is the key towards the succuss of CTR prediction (Rendle, 2010; Wang et al., 2021b). While global features reflect static user interests (Lian et al., 2018; Zhang et al., 2022), the user behavior sequence provides complementary information depicting user interests from a dynamic view (Zhou et al., 2019). For instance, a user’s profile might indicate a general interest in electronics, while the recent browsing history on smartphones offers more specific and timely information about the current need. Therefore, it is crucial to learn behavior-aware global interactions to adapt to the evolving environment.
To learn behavior-aware global interaction, we model the interactions among global features, as well as sequence summarization, by an interaction module. Formally, given the global input $\mathbf{X}^{(l)}$ and sequence summarization $\mathbf{S}_{\text{sum}}^{(l)}$ at the $l$-th layer, the output of $l$-th Global Arch is defined as:
(7) $\mathbf{X}^{(l+1)}=\text{MLP}^{(l)}\left(\text{Interaction}^{(l)}\left(\left[\mathbf{X}^{(l)}\|\mathbf{S}_{\text{sum}}^{(l)}\right]\right)\right),\vspace{-5pt}$

### v3

# InterFormer: Effective Heterogeneous Interaction Learning for Click-Through Rate Prediction

@@ Abstract.
Click-through rate (CTR) prediction, which predicts the probability of a user clicking an ad, is a fundamental task in recommender systems. The emergence of heterogeneous information, such as user profile and behavior sequences, depicts user interests from different aspects. A mutually beneficial integration of heterogeneous information is the cornerstone towards the success of CTR prediction. However, most of the existing methods suffer from two fundamental limitations, including (1) insufficient inter-mode interaction due to the unidirectional information flow between modes, and (2) aggressive information aggregation caused by early summarization, resulting in excessive information loss. To address the above limitations, we propose a novel module named InterFormer to learn heterogeneous information interaction in an interleaving style. To achieve better interaction learning, InterFormer enables bidirectional information flow for mutually beneficial learning across different modes. To avoid aggressive information aggregation, we retain complete information in each data mode and use a separate Cross Arch for effective information selection and summarization. The proposed InterFormer has been deployed across multiple platforms at Meta Ads, achieving 0.15% performance gain and 24% QPS gain compared to prior state-of-the-art models and yielding sizable online impact.
@@ 1. Introduction
Click-through rate (CTR) prediction, which predicts the probability of a user clicking an item, is the fundamental task for various applications such as online advertising and recommender systems (Lyu et al., 2020; Zhang et al., 2022, 2024). The quality of CTR prediction significantly influences company revenue and user experience, drawing extensive attention from academia and industry (Zhou et al., 2018, 2019). For example, in online ad bidding, CTR prediction helps advertisers optimize the bids and target the most receptive audiences. In content recommendation, it enables platforms to suggest relevant content to users.
@@
To achieve better CTR prediction, it is crucial to capture the user interests in the evolving environment (Zhou et al., 2018; Lyu et al., 2020; Wang et al., 2019). The abundance of heterogeneous information presents both opportunities and challenges. On the one hand, heterogeneous information depicts user interests from different aspects, providing diverse context (Zhang et al., 2017). For instance, the non-sequence features, e.g., user profile and context features, offer a static view on general user interests, while behavior sequences provide substantial information for modeling dynamic user interests (Wang et al., 2019). On the other hand, the heterogeneous nature of the data requires different modeling approaches and careful integration across different information modes (Zhang et al., 2017). For example, while modeling interactions among non-sequential information is critical to personalized recommendation (Rendle, 2010; Lian et al., 2018; Wang et al., 2021b), capturing sequential dependencies is the major focus for user behavior modeling (Sun et al., 2019; Chen et al., 2019).
@@
Most of the existing CTR prediction models fall into two categories, including non-sequential and sequential models. Non-sequential models focus on learning informative embeddings through feature interaction via inner-product (Lian et al., 2018; Sun et al., 2021), MLP (Wang et al., 2017, 2021b) and deep structured semantic model (Huang et al., 2013; Elkahky et al., 2015), but ignore the sequential information in user behaviors. Sequential models, in contrast, employ additional modules like CNN (Tang and Wang, 2018), RNN (Sun et al., 2019; Zhou et al., 2018) and Attention modules (Lyu et al., 2020; Zhou et al., 2019; Zhai et al., 2024), to capture the sequential dependencies in user behaviors. Promising as it might be, existing sequential methods mostly employ a unidirectional information flow, where the non-sequential information is used to guide sequence learning, while the reverse information flow from sequence to non-sequence is largely ignored, resulting in insufficient inter-mode interaction. For example, the non-sequential information captures long-term interests, while the sequential information reveals momentary interests, such as a sudden focus on a specific category of products, which can enhance the non-sequential context with immediate preference. Besides, due to the computational challenges of performing interaction learning among numerous non-sequence features and lengthy sequences, aggressive feature aggregation, e.g., sequence summation (Zhou et al., 2018), pooling (Xiao et al., 2020), and concatenation (Zhou et al., 2019), is often performed in the early stages, inevitably leading to excessive information loss.
@@
In light of the above limitations, we propose a novel heterogeneous interaction learning module named InterFormer, whose ideas are two-fold. To avoid insufficient inter-mode interaction, we enable bidirectional information flows between different modes, such that non-sequence and sequence learning are performed in an interleaving style. Specifically, to learn context-aware sequence embeddings, non-sequence summarization guides sequence learning via Personalized FeedForward Network (PFFN) and Multihead Attention (MHA) (Vaswani et al., 2017). To learn behavior-aware non-sequence embeddings, sequence summarization instructs non-sequence learning via an interaction module. To mitigate aggressive information aggregation, we adopt MHA for effective information selection such that the one-to-one mappings between input and output tokens are retained. Note that our framework is compatible with various interaction learning models like DCNv2 (Wang et al., 2021b), DHEN (Zhang et al., 2022), etc.
@@
  • Model Design. A novel heterogeneous interaction learning framework named InterFormer is proposed for effective feature interaction and selective information aggregation. To our best knowledge, we are the first to address the mutual benefits in heterogeneous interaction learning.
  • Experiments and Analysis. Extensive experiments suggest that InterFormer achieves up to 0.14% AUC improvement on benchmark datasets and 0.15% Normalized Entropy (NE) gain on industrial dataset. In addition, the internal deployment at Meta suggests that InterFormer exhibits promising scaling results, achieving 0.15% performance gain and 24% QPS gain compared to SOTA CTR models, and yielding sizable online impact.
@@ 2.2. Sequential Methods
With the recent emergence of sequential information, e.g., user interaction history, in recommender systems, extensive efforts have been made to capture the evolving user interest. A key challenge behind sequential methods is to combine the sequential information with non-sequential information in a mutually beneficial manner. Markov models (Shani et al., 2005; He and McAuley, 2016a; Yang et al., 2020) consider sequential data as a stochastic process over discrete random variables, but the oversimplified Markovian property limits the model capability in capturing long-term sequential dependencies (Quadrana et al., 2018). To model the long-term dependencies, RNN and Attention mechanism are employed as the backbone module for many sequential methods. For example, various attention-based networks are designed (Zhou et al., 2019; Lyu et al., 2020), and Transformer (Devlin et al., 2018; Vaswani et al., 2017) is adopted for sequential modeling (Sun et al., 2019; Chen et al., 2019). Besides, to model multi-faceted user interests, (Xiao et al., 2020; Han et al., 2024) propose to capture multiple user interests from multi-behavior sequences. More recently, TransAct (Xia et al., 2023) adopt a hybrid ranking model to combine real-time user actions for immediate preference and batch user representations for long-term interests. LiRank (Borisyuk et al., 2024) improves CTR prediction at LinkedIn by ensembling multiple interaction modules, accelerated by quantization and vocabulary compression. CARL (Chen et al., 2024) achieves fast inference on large-scale recommendation in Kuaishou by utilizing cached results, monitored by a reinforcement learning-based framework. While most of the existing sequential methods leverage non-sequential information for personalized sequence modeling, sequential information is rarely explored for non-sequence learning. We believe such unidirectional design limits the expressiveness of the learned embeddings, and a bidirectional information flow between different modes is the key towards better heterogeneous interaction learning.
@@ 3. Preliminaries
Table 1 summarizes the main symbols and notations used throughout the paper. We use bold uppercase letters for matrices (e.g., $\mathbf{X}$), bold lowercase letters for vectors (e.g., $\mathbf{x}$), and lowercase letters for scalars (e.g., $n$). The element at the $i$-th row and $j$-th column of a matrix $\mathbf{X}$ is denoted as $\mathbf{X}(i,j)$. The transpose of $\mathbf{X}$ is denoted by the superscript ${\scriptscriptstyle\mathsf{T}}$ (e.g., $\mathbf{X}^{\scriptscriptstyle\mathsf{T}}$). We use superscript $u$ to denote users, and subscripts $i$ and $t$ to denote item and timestamp, respectively (e.g., $y^{u}_{i_{t}}$). We use $\mathbf{x}_{j}^{(l)}$ to denote the $j$-th non-sequence feature of the $l$-th layer, and $\mathbf{s}_{t}^{(l)}$ to denote the sequence feature of the $l$-th layer at timestamp $t$. We consider the scenario with $m$ dense features, $n$ sparse features, and $k$ sequence features of length $T$. We use $d$ to denote the embedding dimension of the CTR model.
@@
$\mathbf{x}_{j}^{(l)}$ $j$-th non-sequence feature of layer $l$
$\mathbf{s}_{t}^{(l)}$ sequence feature of layer $l$ at time step $t$
@@ 3.1. Click-Through Rate (CTR) Prediction
CTR prediction estimates the probability of a user clicking on an item given heterogeneous information such as static context and behavior sequences. Formally, given a user set $\mathcal{U}$ and an item set $\mathcal{I}$, the interaction sequence of a user $u\in\mathcal{U}$ is defined as $S^{u}=[i^{u}_{1},i^{u}_{2},\dots,i^{u}_{T}]$, where $i^{u}_{t}\in\mathcal{I}$ is the item interacted at time step $t$ by user $u$. We aim to estimate the probability of a user clicking on a new item $i^{u}_{T+1}$, denoted as $y^{u}_{i_{T+1}}$, given the historical interaction sequence $S^{u}$. Formally, we seek to learn a function $f:\mathcal{U}\times\mathcal{I}\times\mathcal{S}\rightarrow[0,1]$, where $\mathcal{S}$ is the set of all possible sequences, such that:
@@ 3.2. Feature Interaction
DCNv2 further combines the stacked cross layers for explicit interactions with a deep layer for implicit interactions, enabling an output enriched with comprehensive information.
Deep Hierarchical Ensemble Network (DHEN) (Zhang et al., 2022). DHEN learns a hierarchy of interactions based on a layered structure where each layer is an ensemble of multiple heterogeneous interaction modules. Given the concatenation of $m$ input features $\mathbf{X}^{(l)}\in\mathbb{R}^{d\times m}$, the output of the $l$-th DHEN layer is computed as:
@@ 3.3. Attention Mechanism
Attention mechanisms (Bahdanau et al., 2014) have become an integral part of sequence modeling, serving as the core component behind various model designs (Vaswani et al., 2017; Devlin et al., 2018). We briefly review attention mechanisms, including Multi-Head Attention and Pooling by Multihead Attention (PMA).
@@
(2) $\text{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\mathbf{QK}^{\scriptscriptstyle\mathsf{T}}/\sqrt{d_{k}}\right)\mathbf{V},$
where $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{n\times d_{k}}$ are query, key and value representations computed by linear projections $\mathbf{Q}=\mathbf{S}^{\scriptscriptstyle\mathsf{T}}\mathbf{W}^{Q},\mathbf{K}=\mathbf{S}^{\scriptscriptstyle\mathsf{T}}\mathbf{W}^{K},\mathbf{V}=\mathbf{S}^{\scriptscriptstyle\mathsf{T}}\mathbf{W}^{V}$.
@@
(3) $\text{MHA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\left[\textbf{head}_{1}\|\ldots\|\textbf{head}_{h}\right]\mathbf{W}^{O},$
@@
(4) $\text{PMA}(\mathbf{Q}_{\text{PMA}},\mathbf{S})=\text{MHA}(\mathbf{Q}_{\text{PMA}},\mathbf{K},\mathbf{V}).$
@@ 4. Methodology
Figure 1. An overview of the InterFormer model architecture. Each block consists of three parts, including (1) Interaction Arch (orange) to learn behavior-aware non-sequence embeddings given sequence queries; (2) Sequence Arch (blue) to learn context-aware sequence embeddings given non-sequence queries; (3) Cross Arch (green) to exchange information between Interaction and Sequence Arch. Note that the dashed green line for CLS token appending only happens at the first layer.
In this section, we present our proposed InterFormer. We first introduce the preprocessing module in Section 4.1, followed by three major modules, including Interaction Arch, Sequence Arch and Cross Arch. To learn behavior-aware non-sequence embeddings, the Interaction Arch models feature interactions (Section 4.2). To learn context-aware sequence embeddings, the Sequence Arch is proposed to optimize sequence (Section 4.3). The Cross Arch connects Interaction and Sequence Arches, enabling effective information summarization and exchange between different modes (Section 4.4). An overview of InterFormer is shown in Figure 1.
@@ 4.1. Feature Preprocessing
Non-Sequence Feature Preproessing. Two types of non-sequence features are considered, including dense features like user age and item price, and sparse features like user id and item category. To unify the heterogeneity in non-sequence features, different features are transformed into embeddings of the same dimensionality.
@@
By concatenating the dense and sparse embedding vectors, the input non-sequence embedding matrix can be obtained as follows
@@
Sequence Feature Preprocessing. Similar to the non-sequence feature preprocessing, an embedding layer is employed such that each interacted item in the sequence is mapped to a $d$-dimensional vector $\mathbf{s}_{t}$, and the behavior sequences is represented as the concatenation of item embeddings, i.e., $\mathbf{S}^{(0)}=\left[\mathbf{s}_{1}^{(0)}\|\dots\|\mathbf{s}_{T}^{(0)}\right]\in\mathbb{R}^{d\times T}$.
Real-world scenarios often encounter multiple sequences from different user actions (e.g., click and conversion)or various platforms. Moreover, due to the uncertainty of user behaviors, these sequences often contain noisy and irrelevant user-item interactions. To address this, a MaskNet (Wang et al., 2021a) is employed to unify multiple sequences and filter out the internal noises via self-masking.
@@
where $\text{MLP}_{\text{mask}}\!:\mathbb{R}^{kd\times T}\!\!\to\!\mathbb{R}^{kd\times T}$ computes self-masking to select relevant information from input sequences, and $\text{MLP}_{\text{lce}}:\mathbb{R}^{kd\times T}\!\to\!\mathbb{R}^{d\times T}$ linearly combines multiple sequences into one, matching the dimensionalities of non-sequence and sequence features.
4.2. Interaction Arch: Towards Behavior-Aware Interaction Learning
Non-sequence features, such as user profile and ad content, provide substantial information in understanding user preference over a specific item. Modeling the interaction among non-sequence features is the key towards the success of CTR prediction (Rendle, 2010; Wang et al., 2021b). While non-sequence features reflect static user interests (Lian et al., 2018; Zhang et al., 2022), the behavior sequence provides complementary information depicting user interests from a dynamic view (Zhou et al., 2019). For instance, while user profile exhibits a general interest in electronics, the recent browsing history on smartphones offers more specific and timely information on current needs. Therefore, it is crucial to learn behavior-aware non-sequence interactions to adapt to the evolving environment.
To learn behavior-aware non-sequence interaction, we model the interactions among non-sequence features, as well as sequence summarization, by an interaction module. Formally, given the non-sequence input $\mathbf{X}^{(l)}$ and sequence summarization $\mathbf{S}_{\text{sum}}^{(l)}$ at the $l$-th layer, the output of $l$-th Interaction Arch is defined as:
(7) $\mathbf{X}^{(l+1)}=\text{MLP}^{(l)}\left(\text{Interaction}^{(l)}\left(\left[\mathbf{X}^{(l)}\|\mathbf{S}_{\text{sum}}^{(l)}\right]\right)\right),$